OpenStudy (anonymous):

what is the solution to the system of linear equations 2x+y-6 x-y=0 show your work

OpenStudy (lncognlto):

Is it 2x - y - 6 = 0 or 2x - y = 6?

OpenStudy (anonymous):

2x-y-6

OpenStudy (lncognlto):

Ok then 2x + y - 6 = 0 x - y = 0 then 3x = 6 x=2 then y also = 2.